OpenStudy (anonymous):

7.5 x 10^4 J = 440 g (0.90J) (Tf- 19.5 C) how do I rearrange this to solve for Tf?

OpenStudy (mertsj):

Well first, divide both sides by 440(.90)g-J

OpenStudy (mertsj):

And then add 19.5 to both sides.

OpenStudy (anonymous):

thats it?

OpenStudy (mertsj):

yep

OpenStudy (anonymous):

what about in this case? {155 g}{4.184 J} {Tf – 26C} = - {75 g} {4.184 J} { Tf – 95C}

OpenStudy (mertsj):

Well then you have to use the distributive property and collect the terms that contain Tf

OpenStudy (anonymous):

how?

OpenStudy (mertsj):

\[155(4.184)(T-26)=-75(4.184)(T-95)\]

OpenStudy (mertsj):

Let's divide both sides by 4.184 and be rid of that

OpenStudy (anonymous):

divide or subtract?

OpenStudy (mertsj):

\[155(T-26)=-75(T-95)\] \[155T-4030=-75T+7125\] \[230T=11155\] \[Tf=48.5\]

OpenStudy (mertsj):

Is it a factor or an addend?

OpenStudy (mertsj):

If 4.184 is a factor divide it out. It it is an addend, subtract it. Which is it?

OpenStudy (anonymous):

what calculation did you perform to get the 4030

OpenStudy (anonymous):

and the 7125

OpenStudy (mertsj):

Do you know how to use the distributive property?

OpenStudy (mertsj):

155(T-26)=155 times T - 155 times 26

OpenStudy (anonymous):

a(b+c) = ab + bc

OpenStudy (mertsj):

Yep

OpenStudy (anonymous):

ok i get it now. thanks for you patience

OpenStudy (mertsj):

Also -75(T-95)==-75 time T + 75 times 95

OpenStudy (mertsj):

yw